### v1

# Experimental Analysis of Efficiency of the Messaging Layer Security for Multiple Delivery Services

Abstract
Messaging Layer security (MLS) and its underlying Continuous Group Key Agreement (CGKA) protocol allows a group of users to share a cryptographic secret in a dynamic manner, such that the secret is modified in member insertions and deletions. One of the most relevant contributions of MLS is its efficiency, as its communication cost scales logarithmically with the number of members. However, this claim has only been analysed in theoretical models and thus it is unclear how efficient MLS is in real-world scenarios. Furthermore, practical decisions such as the chosen Delivery Service and paradigm can also influence the efficiency and evolution of an MLS group. In this work we analyse MLS from an empirical viewpoint: we provide real-world measurements for metrics such as commit generation and processing times and message sizes under different conditions. In order to obtain these results we have developed a highly configurable environment for empirical evaluations of MLS through the simulation of MLS clients. Among other findings, our results show that computation costs scale linearly in practical scenarios even in the best-case scenario.
keywords:
Continuous Group Key Agreement , MLS , Delivery Service , Network simulation , Decentralised Communications
journal: Computers & Security
[…]
{highlights}
We study the efficiency of the Messaging Layer Security (MLS) protocol in an experimental setting.
We provide an implementation of a configurable environment for the empirical evaluation of MLS.
We measure latency, generation and processing times and message sizes for different Delivery Services, paradigms and group states.
1 Introduction
Messaging Layer Security (MLS) [1] is a recent communications standard for establishing secure messaging groups between a set of users. It is mainly composed by a cryptographic scheme named Continuous Group Key Agreement (CGKA), whose objective is to distribute a shared secret between members of the group. One of the most relevant characteristics of CGKA protocols is its flexibility in group composition: the scheme allows insertions and deletions of members while ensuring the security properties of Forward Secrecy (FS) and Post-Compromise Security (PCS).
In addition to these security properties CGKA schemes also have a significant focus in maintaining efficiency as the group increases in size. Indeed, MLS employs a version of TreeKEM [2] as its underlying CGKA protocol, which structures the group as a binary tree in order to achieve logarithmic complexity in relation to group size. The efficiency of CGKA protocols has been widely analysed in the literature [3, 4, 5] and multiple CGKA variants that attempt to increase efficiency under certain scenarios have been proposed [6, 7, 8, 9].
However, most contributions to the CGKA literature only address the theoretical aspects of these protocols. While some works perform empirical measurements of their CGKA variants [10, 11, 12], they are often limited to small groups and specific scenarios. This prevents the analysis of practical considerations that also affect the development of MLS groups. Specifically, the Authentication and Delivery services defined in the MLS RFC have received very little attention [13], and are usually instantiated as abstract functionalities. The latter is relevant for the execution of the CGKA protocol, as it handles the transmission of messages between members of a group and stores information required to insert new members. The treatment of these services as abstract and efficient obfuscate their influence on the correct functioning and efficiency of the protocol. Thus, the interaction between the underlying CGKA schemes and the real-world instantiation of these services has not been explored.
Furthermore, MLS introduces new elements to CGKA groups whose impact has not been studied in the literature. In particular, MLS defines a method for External Joins, through which users can enter a group without requiring invitation. The KeyPackage and GroupInfo messages are required to add new members to the group, and they need to be stored and made available to users. The propose-and-commit paradigm requires the generation, distribution and processing of proposal messages that represent an additional overhead.
In this work we address those issues by performing an empirical evaluation of the performance of MLS under multiple different scenarios. We study how performance - mainly commit generation and processing times, as well as message size - is affected by different conditions such as the chosen Delivery Service and paradigm.
To that end we have developed a testbed for experimental evaluations of MLS. Our implementation consists of a simulated MLS client that autonomously joins groups and periodically publishes messages and updates, as well as two different Delivery Services to distribute MLS messages between the clients. The testbed is highly configurable and allows for the deployment of an arbitrary number of simulated clients. We publish our implementation as open-source so that it is available for researchers to perform their own analysis or even create new scenarios.
In summary, our work presents the following contributions:
  1. 1.
    A testbed for experimental evaluations of MLS. We provide a Rust implementation [^1] Available at https://github.com/SDABIS/mls_experimental_analysis of a simulated MLS client that can be configured in detail to perform various tests. When executed in parallel, these clients will create MLS groups, add or remove other clients and exchange application messages, according to their parameters. We also implement two different Delivery Services: a centralised MQTT server and a distributed P2P-based publish-subscribe network.
  2. 2.
    An empirical analysis of the performance of MLS. We study MLS behaviour under multiple conditions: the type of Delivery Service employed, the number and behaviour of users and whether or not some MLS functionalities like External Joins and the propose-and-commit paradigm are being employed. We measure the computational cost of generating and processing updates, the size of exchanged messages and the time required to deliver them to all members. Among other findings, our results show that the expected logarithmic complexity of MLS does not hold in practice, even in a best-case scenario.
The rest of the document is organised as follows: Section 2 will review related works to provide the context to our contribution. Section 3 will introduce to the reader the concepts of CGKA and MLS that are relevant for our analysis. In Section 4 we identify the variables we will measure and the parameters that will impact them. Section 5 introduces our implementation of the simulated MLS clients and the Delivery Services and in Section 4.2 we present our measurements of said environment. In Section 7 we discuss our most relevant findings and overall implications of our results. Finally, Section 8 will conclude this document and discuss future improvements.
2 Related Work
CGKA protocols have been thoroughly defined in the literature [14, 15, 16], exploring the security properties of Forward Security (FS) and Post-Compromise Security (PCS) [14]. Since one of the main objectives of CGKA protocols is to provide efficient scaling with the number of users [5], they usually employ binary trees to represent the state of the group [17, 2, 18]. As such, CGKA protocols usually aim for logarithmic complexity in adds and removals, although the PCS requirement unavoidably causes its efficiency to degrade under certain conditions [19]. The most popular instantiation of a CGKA protocol is the Messaging Layer Security (MLS), which has recently been standardised as RFC 9420 [1]. MLS employs TreeKEM [2] as its CGKA protocol, and uses the shared cryptographic secret to perform Secure Group Messaging between its members.
The efficiency of CGKA protocols has mostly been discussed in a theoretical framework, without presenting experimental results. The authors of [6] discuss the bandwidth required to distribute messages to all members of the CGKA group, and propose introducing a server to more efficiently provide each member with the information they require. A similiar approach is taken in [7], which also optimises the protocol for post-quantum algorithms for public key encryption. The impact of the tree’s state in efficiency is taken into account. Other works further develop this analysis to estimate the communication cost (i.e., the amount of messages required) of healing the group after a compromise. Communication cost is estimated both for a generic CGKA protocol [3, 4, 5] and for specific schemes [20, 21, 22]. The authors of [12] also analyse the communication cost of the TreeKEM protocol by focusing on the shape of the ratchet tree.
The authors of [23] propose a Delivery Service that employs Reliable Broadcast and consensus mechanisms to distribute proposals and commits, respectively. They later presented an implementation [24], but without experimental measurements. The CGKA variant introduced in [8] is particularly suited for its implementation using a Blockchain as a decentralised Delivery Service, although the work is mainly theoretical and no experimental environment is discussed. Conversely, the authors of [25] do implement a CGKA environment using a blockchain for IoT devices, but employ Asynchronous Ratchet Trees [17], an outdated version of TreeKEM and thus not compatible with MLS. In [11] an alternative CGKA protocol that does not employ binary trees is presented and an implementation is provided for a simplified execution environment.
Experimental analysis of CGKA implementations is limited to works that present an specific CGKA variant such as [10, 11, 26], although their experimental settings are limited: they only consider groups with few members - up to 128 - and do not model user behaviour. In both cases their measurements show how commit generation times increase linearly as the number of users grow.
3 Background
3.1 Continuous Group Key Agreement
A Continuous Group Key Agreement (CGKA) protocol is a scheme that allows a set of users to establish a common secret. This shared value changes dynamically with every modification of the state of the group, whether it be insertions or eliminations of users. In addition to the shared secret, each member possesses some private information such as a signing key pair linked to an identity credential. Each state of the group is called an epoch; whenever a member issues a modification to the state of the group through a commit message, a new epoch is created with a different shared secret. Throughout this document, we will refer to the group member who generates a commit as committer.
CGKA protocols are designed to resist the leakage of the shared secret, whether it be because of a failure in the implementation or the compromise of a member, whose secrets are stolen [15]. As a result, the most relevant security properties a CGKA protocol aims to provide are Forward Secrecy (FS) and Post-Compromise Security (PCS) which state that if the shared secret of an epoch is compromised, it should not be possible to obtain any secret information from previous and following epochs.
A commit may add a new member to the group or remove an existing one. Additionally, members can modify their individual state if they believe it has been compromised. This also refreshes the group’s state and may modify some of the group’s contextual information and parameters. We will refer to this actions as Add, Remove and Update, respectively.
Efficiency is also a main focus of CGKA protocols, aiming to achieve logarithmic complexity in updates to the state of the group [5]. To this end, tree structures in which each user possesses information stored in a leaf node [17, 2] are common in the literature.
3.2 Messaging Layer Security (MLS)
MLS uses TreeKEM [2] as its underlying CGKA protocol. The group’s state is structured as a binary tree in which every node holds a cryptographic key pair. The leaf nodes represent the members and contain other information like their credentials and signature key. The participants only know the secret key of those nodes that are included in their path to the root. The secret contained in the root node is known to all members, and thus it is employed to derive the shared secret.
Commits that alter the group’s state may modify both the committer’s leaf node and the secret key of intermediate nodes in its path to the root. In order to securely transmit these changes to other users, a set of path secrets is generated by encrypting the new secrets with some of the tree’s public keys, such that all members can recalculate the shared secret.
For the correct functioning of an MLS group, two abstract services are defined [27]: the Authentication Service (AS) and the Delivery Service (DS). The former is tasked with generating the member’s credentials and assisting in validating the identity of other members. The AS is considered to be an external entity to the protocol: for example, it could be instantiated as the current Public Key Infrastructure, in which X.509 certificates containing a public key are signed by Certificate Authorities that are trusted by default.
Figure 1: Message flow in MLS through the Delivery Service.
On the other hand, the Delivery Service is tasked with distributing messages and storing relevant information about both groups and users. The flow of information in MLS is shown in Figure 1: Whenever group members send a message, it is published to the Delivery Service, which is tasked with forwarding it to the recipients (usually the other group members).
3.3 Communication in MLS
The following message types are defined in MLS [1]:
  • 1.
    Handshake: Messages that apply or propose changes to the group’s state. Usually handshake messages are generated by group members, with the exception of External Joins, in which non-members insert themselves into the group. By default, handshake messages are encrypted with the group’s shared secret. Handshake messages fall into the following categories:
    • (a)
      Proposal: does not apply a change directly to the group, but instead communicates the intention to do so.
    • (b)
      Commit: contains one or more proposals (not necessarily generated by the committer) and applies them to the group. It also contains the updated path secrets that the receivers will employ to update their Ratchet Tree.
  • 2.
    Welcome: Generated by commits that include new members to the group. It contains a copy of the group’s Ratchet Tree and any other information that the new user needs in order to participate in the group.
  • 3.
    Application: any message exchanged between group members, protected by the group’s secret.
These messages are generated by users and distributed to members of the group (in the case of the Welcome message, to the new member) through the Delivery Service.
In addition to delivering protocol messages, the DS must also store some information generated by users and groups such that it can be accessed by any party on demand. We will refer to this abstract functionality of the DS as Directory. There are two types of objects that the Directory must store:
  • 1.
    Key Packages: they are generated by users and contain cryptographic information required to insert them into an MLS group, including their credential and capabilities as well as an Hybrid Public Key Encryption (HPKE) [28] public key. Key Packages are included in Add proposals and cannot be reused.
  • 2.
    Group Information (GroupInfo): they contain the a group’s Ratchet tree and other contextual information that is required for performing External Joins. It also contains an HPKE public key whose secret counterpart is derived from the group’s shared secret, and thus is known by all members. GroupInfo packages can be generated by any group member and represent its state for an specific epoch.
Since MLS messages are encrypted, the DS does not need to be a trusted party. However, we remark that a malicious or compromised DS still has the capability of tracking group compositions, reordering or deleting messages or even deny service to a party.
3.4 Group Updates
(a) Invitation.
(b) External Join.
Figure 2: Overview of the Invitation and External Join methods through which new members are added to the MLS group.
As mentioned, a group’s state is modified by commits which can either add or remove members or update their individual state. While the two latter operations are fairly simple, inserting new members into an MLS group requires additional steps. They can either be added through Invitation - if a current member inserts them into the group - or External Join - if they insert themselves [13]. These two methods are represented in Figure 2.
Invitation
Figure 2(a) shows the message flow for inviting new users to the group. As a prerequisite, the new user Bob must have uploaded a Key Package created by him to the Delivery Service’s Directory. Then, the group member member Alice obtains it and generates an Add proposal that includes the Key Package. When the proposal is committed, Alice employs the HPKE public key contained in Bob’s Key Package to encrypt the group’s secrets and includes them in a Welcome message directed to Bob, along other contextual information required to participate in the group.
All other members of the group then process Alice’s commit. Bob’s Key Package also included a leaf node which is inserted into the group’s TreeKEM after validating his credentials.
External Join
In contrast to Add proposals, External Joins are initiated by the joiner, who is not a member of the group. The procedure is shown in Figure 2(b): First, the joiner Bob accesses the group’s GroupInfo, which was uploaded to the Directory by the group member Alice. Then, he generates a value that will be used to derive the new epoch’s shared secret and encrypts it with the group’s HPKE public key (contained in the GroupInfo) such that it is accessible only to members of the group. Finally, Bob generates a commit with the Join proposal - a new proposal type defined by MLS -, which is validated and processed by all members of the group. We remark that an MLS group can specify if External Joins are allowed.
4 Empirical evaluation of MLS
In this Section we will analyse the most relevant properties of the execution environment in which we will measure the behaviour of MLS using different Delivery Services. We will start by defining the parameters that will configure our testbed. Then, we will introduce the variables that we will measure, whose values will be predictably affected by the parameter configuration.
4.1 Parameters
The most relevant metric for our experiments is group size, that is, the current number of members of the MLS group. It is claimed that MLS’ efficiency scales logarithmically with group size [6, 20], and all of our measurements are is significantly influenced by it. We will also analyse how the other parameters introduced in this Section interact with group size.
In addition to group size, we will measure the influence of the following parameters:
4.1.1 Delivery Service
The Delivery Service can be instantiated as a centralised server that manages the MLS groups and relays messages to its participants. The use of a server help to ensure consistency as it can ensure that all parties receive all messages and in the same order. However, the server also represents a single point of failure and may introduce bandwidth limitations, as all messages need to go through a single machine. On the other hand, a distributed network of devices that autonomously transmit and forward messages between them can also serve as a Delivery Service. In addition to the distribution of messages, the distributed network can also serve as Directory by storing KeyPackages and GroupInfo messages.
4.1.2 Paradigm
Users apply modifications to an MLS group by issuing a commit, which updates the shared secret and certain nodes in the Ratchet Tree. This can lead to inefficiencies if multiple commits are generated consecutively, as the updated secrets are needlessly overwritten. Alternatively, a commit can simultaneously apply multiple modifications through the use of proposals. We refer to this two different approaches as paradigms.
More specifically, we will distinguish between these paradigms:
  • 1.
    Commit: Modifications are directly and individually applied to the group by a Commit.
  • 2.
    Propose-and-commit: A single commit includes and applies any number of previously exchanged proposals, possibly generated by different members. This effectively reduces the amount of commits generated while maintaining the overall behaviour of the group.
The propose-and-commit paradigm reduces the number of modifications to the Ratchet Tree that are required to perform the same functionality. However, it also introduces an additional overhead: the generation, distribution and processing of proposal messages. In the propose-and-commit paradigm, a commit that applies $n$ proposals requires the exchange of $n+1$ messages: each of the $n$ proposal messages and the commit message itself. This contrasts with the commit paradigm, in which every handshake message applies exactly one modification. We unify these results under the following Average Update cost ($auc$) metric:
$auc=\frac{cc+\sum_{i=0}^{n}cp_{i}}{n}$
where $cc$ and $cp_{i}$ are the costs of the commit and its $i$th proposal, respectively.
We will also consider External Joins in our discussion of paradigms. Although External Joins are not incompatible with any of the other two paradigms - as removes and updates can still be either commit or propose-and-commit - they do represent a different method for adding new members to the group that can be compared with them.
4.1.3 Ratchet Tree State
Although the TreeKEM protocol aims to achieve to scale logarithmically as the number of users grow, this is only the ideal case: when all the tree branches, included intermediate nodes, are populated. These branches can be blanked by some operations, such as invitations of removals under certain conditions [18]. In the worst case (i.e. whenever all intermediate nodes are blanked), the group is structured as a list, which implies that the scaling becomes linear [15].
The amount and location of blanked nodes depends mostly on which users perform the updates. A group reaches its worst possible state if all updates are issued by the same user - the group creator -, as only the nodes in its path to the root will be populated. In contrast, the best state is reached when all members issue an update, without considering removes. This ensures that all intermediate nodes are populated. This can be achieved by only allowing the last user to join to issue updates.
We will model these scenarios by restricting who is able to issue updates: the First user for the worst state and the Last for the best case. We will compare them to a normal scenario in which updates are performed by Random users.
4.2 Measurements
We will measure the following values in our executions:
Latency
In the context of this work, we define an user’s latency with respect to a commit as ”the amount of time between the generation of the commit and the moment in which its changes are applied to the user”. Since an MLS group is composed of multiple users, we will measure the average latency and max latency for each commit generated. The latter represents the delay between the generation of a commit and the start of the new epoch for all members.
The time between the generation of a message and its processing is relevant for all messaging applications, but is particularly relevant for handshake messages in MLS: when a member issues a commit, the group remains in an unstable state until all parties have processed the change and thus cannot adequately exchange any other handshake or application messages.
Commit generation and processing times
CGKA protocols that employ tree structures, such as MLS, are claimed to have logarithmic complexity in relation to the number of members [6, 20]. These claims refer to the Communication cost, that is, the amount of information that needs to be exchanged in order to heal the group. This metric’s relation to computational costs is unclear, and thus in this work we analyse both generation and processing time. We will use CPU time to measure both values.
Besides group size, we measure how the choice of paradigm and state of the Ratchet Tree affect generation and processing times.
Message size
Our analysis includes handshake messages - both commits and proposals, using the Average Update cost metric - as well as other message types like Welcome and GroupInfo packages. The latter two include the full Ratchet Tree, so their size is expected to scale poorly as the group grows.
This experimental measurement complements the communication cost that has been widely analysed in the literature. Our measurements can provide specific values that can be fed into communication cost calculations such that the obtained results can be applied to real-world scenarios.
5 Implementation
In this Section we will introduce our implementation of the different Delivery Services we will employ in our execution environment. We have developed our implementation in Rust, using the OpenMLS project [29] as a baseline and incrementally expanding upon it.
We employ two different Delivery Services, each of them with their corresponding Directory. The first employs a centralised server and forwards messages through the MQTT protocol while the second distributes them through a peer-to-peer network. Their implementations are analysed in subsequent sections. Both are based in a Publish-Subscribe architecture.
Figure 3 shows the handling of topics of a Publish/Subscribe Delivery Service for MLS. Every group has a different topic to which all current members are subscribed. Every user is subscribed to a welcome topic through which it receives all Welcome messages directed to them. Additionally, all groups have their own topic that all members are subscribed to, and every handshake and application message of that group is published through it.
Figure 3: Subscription process for new group members.
5.1 MQTT Delivery Service
MQTT is a protocol for establishing publish/subscribe queues of messages [30]. In this architecture, users can publish messages to different topics such that they are received all users that are subscriber to said topic. The transmission of these messages to all subscribers is managed by a centralised server named MQTT broker. For our execution environment, we employ the open-source Mosquitto [31] implementation of a MQTT broker.
Besides the MQTT broker, the Delivery Service also employs a web server that acts as Directory, with endpoints to store and consume KeyPackages and GroupInfo messages. The latter needs to be updated for each new group epoch, so users aiming to perform an External Join can successfully access the group. Furthermore, this web server also acts as a signaling server in which users can register to notify other participants of their existence, such that they can be invited to MLS groups through an Add proposal.
5.2 GossipSub Delivery Service
For this Delivery Service we maintain the Publish-Subscribe nature but replace the centralised MQTT server with a distributed architecture. We employ the GossipSub protocol [32] that allows peers to subscribe to different topics; peers share with each other metadata of messages they have received and can request the full message if it is from a topic they are subscribed to. Connections are dynamic and favour other members subscribed to the same topic.
We also employ a Decentralised Hash Table (DHT) as this Delivery Service’s Directory. Each client publishes its KeyPackage as a key-value pair, with the key referencing its user identifier. We limit the involvement of the server to that of signaling server as defined for the MQTT Delivery Service.
For our implementation we include Rust’s Libp2p [33] into the simulated clients, which provides an implementation for GossipSub and the Kademlia [34] DHT.
5.3 Client Simulation
Our testbed is composed of multiple simulated clients that independently participate in MLS groups. They emulate the behaviour of real clients by randomly adding or removing members, updating their individual leaf nodes and sending application messages.
[cgka] ds = "mqtt" groups = ["group_1", "group_2"] external_join = true join_chance = 0.01 issue_update_chance = 0.2 message_chance = 0.3 scale = false auth_policy = "Random" message_length_min = 500 message_length_max = 2000 sleep_millis_min = 20000 sleep_millis_max = 60000 [paradigm] paradigm = "propose" proposals_per_commit = 4 invite_chance = 0.6 remove_chance = 0.1 update_chance = 0.3 [http_server] url = "http://<ip>:<port>" [mqtt] url = "tcp://<ip>:<port>" [meta] replicas = 10
Figure 4: Example Client configuration file.
The behaviour of our simulated clients can be specified in detail through configuration files, as shown in Figure 4. Among other parameters, it is possible to configure the groups the user will try to join (or create if they do not exist), how often it will attempt to perform an action and with what probability. The configuration file also defines which paradigm to use and the types of actions (Add, Remove, Update, Join) the client is execute, as well as the Delivery Service that will be employed.
group_1 8 User_0d4341e0c0f4 Invite User_b3f20ed42c2a 1065 1739176120380282661 5885 group_1 8 User_d13041578e84 Process User_0d4341e0c0f4 1739176120384444036 8599 group_1 8 User_b3f20ed42c2a Welcome User_0d4341e0c0f4 2755 1739176120461981624 1823
Figure 5: Logs generated by clients.
Every action related to the evolution of an MLS group is recorded in a log file. Figure 5 shows the trace of an Invitation, which includes involves the following entries: Invite, Process and Welcome. Each entry records the size of the message, the timestamp at which it was performed and its time cost.
6 Experimental Results
In this section we deploy the environment presented in previous sections and perform empirical evaluations on its performance.
We have executed all following tests three times with the objective of clearing any potential noise in the measurements. Every figure in this Section shows the average of the results obtained in all executions; we omit the standard deviation as it is not relevant to our scenarios.
6.1 Latency
(a) MQTT
(b) GossipSub
Figure 6: Mean and Max Latency for the Delivery Services.
Figures 6(a) and 6(b) show the latency for the two Delivery Services. While both experience a significant increase as the number of users grow, the GossipSub DS is noticeably less efficient as its latency more than doubles that of the MQTT DS. Latency also increases more irregularly of GossipSub, with an important increase at 512 group members. However, at higher group sizes the latency grows slower for GossipSub.
6.2 Paradigm
We now compare the Average Update cost for the commit and propose-and-commit paradigms, including multiple amounts of proposals per commit for the latter to see how the cost varies as the number of updates increases. We note that our measurements do not include the cost of retrieving information from the Directory such as KeyPackages for Add proposals, as it remains constant regardless of the chosen paradigm.
(a) Commit and propose-and-commit paradigms
(b) Commits (also including Welcome processing time) and External Joins.
Figure 7: Average Update generation time for different paradigms as the number of users grow.
(a) Processing time.
(b) Message size.
Figure 8: Average Update cost for different paradigms as the number of users grow.
Figures 7(a) and 7(b) show the Average Update generation time for the two paradigms and for External Joins, respectively. Clearly, the Average Update cost is significantly reduced as the number of proposals increases.
We also note that the time to perform an External Join is much larger to that of inviting new members through Add proposals. In order to understand these differences we highlight that an user performing an External Join must also parse the current state of the group including the full Ratchet Tree, which would explain the poor scaling of External Joins. This operation is also performed in an Invitation setting by the invited user when the Welcome message is processed; when accounting for the processing time of Welcome messages, the two methods have a similar cost (as shown in Figure 7(b) by the line labelled Commit + Welcome).
Figure 8(a) shows the processing time per Update as the number of users grow. As with generation time, the processing cost is significantly reduced as the number of proposals increases. Processing time is noticeably more stable than generation time, although it experiences significant increases at the powers of 2 as a new layer is added to the Ratchet Tree. It is apparent that both generation and processing times scale mostly linearly with the number of members in the CGKA group, with $R^{2}$ scores over 0.9 when adjusting for linear regression and around 0.6 for logarithmic regression.
Similarly, Figure 8(b) shows the message size per update. Clearly, the correlation between the message size and the number of users is much smaller than in the other metrics. Contrary to the results obtained while measuring processing and generation times, message size is generally bigger as the number of proposals increase. We recall that commit messages also include copies of the proposals it applies; this redundant information increases the cost per update.
6.3 Other Messages
Figure 9: Message size for Welcome messages and GroupInfo packages.
We also evaluate the size of the Welcome messages and GroupInfo packages, which take part in Invitations and External Joins, respectively. Figure 9 shows the obtained results. Since both contain a full copy of the Ratchet Tree, they are significantly larger than any handshake message. Expectedly, their size increases linearly with the number of group members.
6.4 Ratchet Tree State
(a) Generation time.
(b) Message size.
Figure 10: Generation time and message size for different states of the group’s Ratchet Tree. Processing time is omitted as it is not influenced by this metric.
Finally, we analyse the impact in performance of the current state if the Ratchet Tree, exemplified by the First, Last and Random scenarios introduced in Section 4.1. Figure 10 shows the generation time and message size - processing time is omitted as no significant difference was found between the three scenarios.
As with previous measurements, generation time also scales linearly in all scenarios - this includes Last which corresponds to the best possible state of the tree. Message size does show a significant correlation, with First increasing much faster than in the other scenarios.
7 Discussion
Comparison with other works
Protocol Perspective Group Size Time Bandwidth Tree State Paradigm Join DS
[6] CGKA Variant Theoretical N/A
[3, 4] CGKA Theoretical N/A
[11] CGKA Variant Experimental $\approx$ 128
[10] CGKA Variant Experimental $\approx$ 128
[24] MLS Theoretical N/A
Ours MLS Experimental $\approx$ 1000
Table 1: Comparison between works that analyse CGKA performance and ours. The columns indicate if the works consider the following topics. CPU: update generation/processing time. Bandwidth: message size and communication cost. Paradigm: influence of chosen paradigm. Tree State: Influence of the current state of the Ratchet Tree. Join: External Joins. DS: influence of the Delivery Service.
Table 1 shows a comparison between our work and other works that perform an analysis of performance of CGKA. Our experimental approach allows us to measure CPU time in generating and processing messages and the importance of the Delivery Service in their distribution. When compared with other experimental works, ours has the largest group size and also covers less studied topics such as the influence of paradigm and the current state of the Ratchet Tree.
Delivery Service
The increased bandwidth of larger groups has a noticeable impact on latency, as the cost of distributing messages between all members increases. The MQTT DS performs significantly better than the GossipSub DS, but the former does not require a centralised server to distribute the messages and thus is more reliable.
Latency represents the most significant cost in the evolution of MLS groups, as every other generation or processing cost is negligible in comparison. Updates in MLS are necessarily sequential: if two commits are generated from the same epoch, the group enters into an inconsistent state. In order to generate a valid commit, an user must have received and processed all previous valid commits. Thus, latency is a significant bottleneck to group evolution as it represents the minimum amount of time between two valid commits.
Our measurements reach up to 2 seconds of maximum latency in which the group is in an inconsistent state; while for most applications this span is too short to be noticed, it is something that must be taken into account for situations that require strict synchronisation.
Message cost
Generation and processing times are comparable and represent a small cost in the order of milliseconds. Both experience noticeable increases when a new layer is added to the Ratchet Tree, that is, when group size reaches the next power of 2. In practice, both generation and processing times increase linearly with the amount of users in the group, which contradicts the theoretical assumptions that the protocol complexity is logarithmic. Notably, this trend is maintained even in the best-case scenario where the minimum amount of intermediate nodes are blanked. When compared to latency, with generation and processing times represent a negligible cost.
Welcome messages and GroupInfo packages are significantly costlier to process - as shown in Figure 7(b) - than normal commits. Fortunately, they only need to be processed once in order to access a group. This cost difference is due to their size as they both contain the full Ratchet Tree, which needs to be parsed by the receiving party. In contrast, the size of handshake messages is less influenced by group size and is mostly determined by the state of the Ratchet Tree. Generation time is also significantly affected. Even though these two metrics are slightly correlated they present significant differences and thus should be analysed separately. Our results show that groups in which only a small number of users generate updates are less efficient.
Impact of paradigm
The propose-and-commit paradigm also significantly affects performance: the cost of generating and processing proposal messages is negligible compared to the cost of applying an update to the group. As shown, more proposals per update increase the efficiency of the protocol, but we remark that such amount of modifications to a group applied at once would be rare in a real setting, only possible in a highly unstable group. The size of handshake messages is less affected by the amount of group members and the paradigm, as it is mostly determined by the current state of the Ratchet Tree.
External Joins combine the operations of generating a commit and processing a Welcome message, in the sense that the joiner needs to parse the full Ratchet Tree in order to join. When taking these two operations into account, the cost of an Invitation is similar to that of an External Join.
8 Conclusion and Future Work
In this work we have analysed the performance of MLS in an empirical setting. We analyse the protocol under different parameters such as Delivery Service, chosen paradigm or state of the Ratchet Tree. Our results show that theoretical claims of logarithmic complexity in communication cost do not necessarily manifest in practice.
Our work also demonstrates the relevance of practical considerations in the performance of MLS, as opposed to theoretical analysis. We have found that the most significant time cost - by multiple orders of magnitude - is message distribution through the Delivery Service. Furthermore, the messages that introduce higher workload are Welcome and GroupInfo messages, as they contain the full Ratchet Tree. To the extent of our knowledge, no other work has attempted to increase the efficiency of these operations.
We also publish our execution environment as open source, which includes a configurable simulated MLS client and two different Delivery Services. We hope that it can help other researchers to perform their own simulations and that it can serve as a baseline for more complex simulation environments.
We plan on continuing to study the applicability of concepts defined specifically in MLS - such as the propose-and-commit paradigm or External Joins [13] - to general definitions of CGKA and their impact on the security properties and efficiency of these schemes.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgement
D.S. acknowledges support from Xunta de Galicia and the European Union (European Social Fund - ESF) scholarship [ED481A-2023-219].
This work is part of the project TED2021-130369B-C31 and TED2021-130492B-C21 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”.

### v2

@@
-Experimental Analysis of Efficiency of the Messaging Layer Security for Multiple Delivery Services
+arXiv:2502.18303v2[cs.CR] 27 Feb 2026
 Abstract
@@
 However, this claim has only been analysed in theoretical models and thus it is unclear how efficient MLS is in real-world scenarios.
-Furthermore, practical decisions such as the chosen Delivery Service and paradigm can also influence the efficiency and evolution of an MLS group.
+Furthermore, practical considerations such as the chosen paradigm and the evolution of the group can also influence the performance of an MLS group.
 In this work we analyse MLS from an empirical viewpoint: we provide real-world measurements for metrics such as commit generation and processing times and message sizes under different conditions.
-In order to obtain these results we have developed a highly configurable environment for empirical evaluations of MLS through the simulation of MLS clients.
-Among other findings, our results show that computation costs scale linearly in practical scenarios even in the best-case scenario.
+In order to obtain these results we have developed a highly configurable environment for empirical evaluations of MLS through the emulation of MLS clients.
+Among other findings, our results show that computation costs scale linearly in practical settings even in the best-case scenario.
 keywords:
-Continuous Group Key Agreement , MLS , Delivery Service , Network simulation , Decentralised Communications
-journal: Computers & Security
-[…]
-[…]
-{highlights}
-We study the efficiency of the Messaging Layer Security (MLS) protocol in an experimental setting.
-We provide an implementation of a configurable environment for the empirical evaluation of MLS.
-We measure latency, generation and processing times and message sizes for different Delivery Services, paradigms and group states.
+Continuous Group Key Agreement , MLS , Delivery Service , Performance Evaluation , MQTT
 1 Introduction
 Messaging Layer Security (MLS) [12] is a recent communications standard for establishing secure messaging groups between a set of users.
-It is mainly composed by a cryptographic scheme named Continuous Group Key Agreement (CGKA), whose objective is to distribute a shared secret between members of the group.
+It is mainly composed of a cryptographic scheme named Continuous Group Key Agreement (CGKA), whose objective is to distribute a shared secret between members of the group.
 One of the most relevant characteristics of CGKA protocols is its flexibility in group composition: the scheme allows insertions and deletions of members while ensuring the security properties of Forward Secrecy (FS) and Post-Compromise Security (PCS).
@@
 However, most contributions to the CGKA literature only address the theoretical aspects of these protocols.
-While some works perform empirical measurements of their CGKA variants [10, 11, 12], they are often limited to small groups and specific scenarios.
-This prevents the analysis of practical considerations that also affect the development of MLS groups.
-Specifically, the Authentication and Delivery services defined in the MLS RFC have received very little attention [13], and are usually instantiated as abstract functionalities.
-The latter is relevant for the execution of the CGKA protocol, as it handles the transmission of messages between members of a group and stores information required to insert new members.
-The treatment of these services as abstract and efficient obfuscate their influence on the correct functioning and efficiency of the protocol.
-Thus, the interaction between the underlying CGKA schemes and the real-world instantiation of these services has not been explored.
+Although some works perform empirical measurements of their CGKA variants [11, 34, 17], they are often limited to small groups and specific scenarios.
 Furthermore, MLS introduces new elements to CGKA groups whose impact has not been studied in the literature.
@@
 In this work we address those issues by performing an empirical evaluation of the performance of MLS under multiple different scenarios.
-We study how performance - mainly commit generation and processing times, as well as message size - is affected by different conditions such as the chosen Delivery Service and paradigm.
+We analyse how measurements such as commit generation, processing and distribution times or message size are affected by different experimental conditions: the size of the group, the chosen paradigm, or the evolution of the group state.
 To that end we have developed a testbed for experimental evaluations of MLS.
-Our implementation consists of a simulated MLS client that autonomously joins groups and periodically publishes messages and updates, as well as two different Delivery Services to distribute MLS messages between the clients.
-The testbed is highly configurable and allows for the deployment of an arbitrary number of simulated clients.
+Our implementation consists of a emulated MLS client that autonomously joins groups and periodically publishes messages and updates, as well as two different Delivery Services to distribute MLS messages between the clients.
+The testbed is highly configurable and allows for the deployment of an arbitrary number of emulated clients.
 We publish our implementation as open-source so that it is available for researchers to perform their own analysis or even create new scenarios.
@@
     A testbed for experimental evaluations of MLS.
-We provide a Rust implementation [^1] Available at https://github.com/SDABIS/mls_experimental_analysis of a simulated MLS client that can be configured in detail to perform various tests.
+We provide a Rust implementation [^1] Available at https://github.com/SDABIS/mls_experimental_analysis of a emulated MLS client that can be configured in detail to adapt their behaviour.
 When executed in parallel, these clients will create MLS groups, add or remove other clients and exchange application messages, according to their parameters.
-We also implement two different Delivery Services: a centralised MQTT server and a distributed P2P-based publish-subscribe network.
   2. 2.
     An empirical analysis of the performance of MLS.
-We study MLS behaviour under multiple conditions: the type of Delivery Service employed, the number and behaviour of users and whether or not some MLS functionalities like External Joins and the propose-and-commit paradigm are being employed.
+We define various experiments in which the emulated clients are configured to recreate MLS groups with different configurations.
+The experiments performed study the impact of MLS functionalities like External Joins and the propose-and-commit, as well as the degradation of the group’s state by blanked intermediate nodes.
 We measure the computational cost of generating and processing updates, the size of exchanged messages and the time required to deliver them to all members.
-Among other findings, our results show that the expected logarithmic complexity of MLS does not hold in practice, even in a best-case scenario.
+Among other findings, our results show the relevance of practical considerations such as the chosen paradigm and persistent storage.
+Furthermore, we identify that the expected logarithmic complexity of MLS does not hold in practice even in a best-case scenario due to the high cost of managing the group state, whose size increases linearly as the number of members grow.
 The rest of the document is organised as follows: Section 2 will review related works to provide the context to our contribution.
 Section 3 will introduce to the reader the concepts of CGKA and MLS that are relevant for our analysis.
-In Section 4 we identify the variables we will measure and the parameters that will impact them.
-Section 5 introduces our implementation of the simulated MLS clients and the Delivery Services and in Section 4.2 we present our measurements of said environment.
+In Section 4 we describe our testbed and the experiments to be performed.
+Section 5 introduces our implementation of the emulated MLS clients and the Delivery Service and in Section 6 we present our measurements of said environment.
 In Section 7 we discuss our most relevant findings and overall implications of our results.
@@
 2 Related Work
-CGKA protocols have been thoroughly defined in the literature [14, 15, 16], exploring the security properties of Forward Security (FS) and Post-Compromise Security (PCS) [14].
+CGKA protocols have been thoroughly defined in the literature [5, 8, 6], exploring the security properties of Forward Security (FS) and Post-Compromise Security (PCS) [5, 15].
 Since one of the main objectives of CGKA protocols is to provide efficient scaling with the number of users [16], they usually employ binary trees to represent the state of the group [20, 14, 25].
-As such, CGKA protocols usually aim for logarithmic complexity in adds and removals, although the PCS requirement unavoidably causes its efficiency to degrade under certain conditions [19].
 The most popular instantiation of a CGKA protocol is the Messaging Layer Security (MLS), which has recently been standardised as RFC 9420 [12].
-MLS employs TreeKEM [2] as its CGKA protocol, and uses the shared cryptographic secret to perform Secure Group Messaging between its members.
 The efficiency of CGKA protocols has mostly been discussed in a theoretical framework, without presenting experimental results.
 The authors of [7] discuss the bandwidth required to distribute messages to all members of the CGKA group, and propose introducing a server to more efficiently provide each member with the information they require.
-A similiar approach is taken in [7], which also optimises the protocol for post-quantum algorithms for public key encryption.
+A similar approach is taken in [22], which also optimises the protocol for post-quantum algorithms for public key encryption.
 The impact of the tree’s state in efficiency is taken into account.
@@
 The CGKA variant introduced in [3] is particularly suited for its implementation using a Blockchain as a decentralised Delivery Service, although the work is mainly theoretical and no experimental environment is discussed.
-Conversely, the authors of [25] do implement a CGKA environment using a blockchain for IoT devices, but employ Asynchronous Ratchet Trees [17], an outdated version of TreeKEM and thus not compatible with MLS.
+Conversely, the authors of [1] do implement a CGKA environment using a blockchain for IoT devices, but employ ART [20], an outdated version of TreeKEM and thus not compatible with MLS.
 In [34] an alternative CGKA protocol that does not employ binary trees is presented and an implementation is provided for a simplified execution environment.
-Experimental analysis of CGKA implementations is limited to works that present an specific CGKA variant such as [10, 11, 26], although their experimental settings are limited: they only consider groups with few members - up to 128 - and do not model user behaviour.
+Experimental analysis of CGKA implementations is limited to works that present an specific CGKA variant such as [11, 34, 33], although their experimental settings are limited: they only consider groups with few members —up to 128— and do not model user behaviour.
 In both cases their measurements show how commit generation times increase linearly as the number of users grow.
@@
 Throughout this document, we will refer to the group member who generates a commit as committer.
-CGKA protocols are designed to resist the leakage of the shared secret, whether it be because of a failure in the implementation or the compromise of a member, whose secrets are stolen [15].
-As a result, the most relevant security properties a CGKA protocol aims to provide are Forward Secrecy (FS) and Post-Compromise Security (PCS) which state that if the shared secret of an epoch is compromised, it should not be possible to obtain any secret information from previous and following epochs.
+CGKA protocols are designed to resist the leakage of the shared secret by providing both Forward Secrecy (FS) and Post-Compromise Security (PCS) across epochs.
 A commit may add a new member to the group or remove an existing one.
@@
 This also refreshes the group’s state and may modify some of the group’s contextual information and parameters.
-We will refer to this actions as Add, Remove and Update, respectively.
 Efficiency is also a main focus of CGKA protocols, aiming to achieve logarithmic complexity in updates to the state of the group [16].
@@
 The former is tasked with generating the member’s credentials and assisting in validating the identity of other members.
-The AS is considered to be an external entity to the protocol: for example, it could be instantiated as the current Public Key Infrastructure, in which X.509 certificates containing a public key are signed by Certificate Authorities that are trusted by default.
-Figure 1: Message flow in MLS through the Delivery Service.
-On the other hand, the Delivery Service is tasked with distributing messages and storing relevant information about both groups and users.
-The flow of information in MLS is shown in Figure 1: Whenever group members send a message, it is published to the Delivery Service, which is tasked with forwarding it to the recipients (usually the other group members).
+The Delivery Service is tasked with distributing messages and storing relevant information about both groups and users.
 3.3 Communication in MLS
@@
     Handshake: Messages that apply or propose changes to the group’s state.
-Usually handshake messages are generated by group members, with the exception of External Joins, in which non-members insert themselves into the group.
-By default, handshake messages are encrypted with the group’s shared secret.
-Handshake messages fall into the following categories:
-    • (a)
-      Proposal: does not apply a change directly to the group, but instead communicates the intention to do so.
-    • (b)
-      Commit: contains one or more proposals (not necessarily generated by the committer) and applies them to the group.
-It also contains the updated path secrets that the receivers will employ to update their Ratchet Tree.
+Handshake messages can either be Proposals that communicate the intention to apply a change to the group but do not directly apply it; or Commits that contain one or more proposals and apply them to the group.
   • 2.
@@
 We will refer to this abstract functionality of the DS as Directory.
-There are two types of objects that the Directory must store:
-  • 1.
-    Key Packages: they are generated by users and contain cryptographic information required to insert them into an MLS group, including their credential and capabilities as well as an Hybrid Public Key Encryption (HPKE) [28] public key.
-Key Packages are included in Add proposals and cannot be reused.
-  • 2.
-    Group Information (GroupInfo): they contain the a group’s Ratchet tree and other contextual information that is required for performing External Joins.
-It also contains an HPKE public key whose secret counterpart is derived from the group’s shared secret, and thus is known by all members.
-GroupInfo packages can be generated by any group member and represent its state for an specific epoch.
-Since MLS messages are encrypted, the DS does not need to be a trusted party.
-However, we remark that a malicious or compromised DS still has the capability of tracking group compositions, reordering or deleting messages or even deny service to a party.
+The Directory stores KeyPackages, which are generated by users and contain cryptographic information required to insert them into an MLS group.
+They are included in Add proposals.
+Group Information (GroupInfo) packages are also stored in the Directory, and they contain the group’s Ratchet tree and other contextual information that is required for performing External Joins.
 3.4 Group Updates
@@
 As mentioned, a group’s state is modified by commits which can either add or remove members or update their individual state.
-While the two latter operations are fairly simple, inserting new members into an MLS group requires additional steps.
-They can either be added through Invitation - if a current member inserts them into the group - or External Join - if they insert themselves [13].
-These two methods are represented in Figure 2.
-Invitation
+Although the two latter operations are fairly simple, inserting new members into an MLS group requires additional steps.
+They can either be added through Invitation —if a current member inserts them into the group— or External Join —if they insert themselves [21].
 Figure 1(a) shows the message flow for inviting new users to the group.
@@
 Then, the group member member Alice obtains it and generates an Add proposal that includes the Key Package.
-When the proposal is committed, Alice employs the HPKE public key contained in Bob’s Key Package to encrypt the group’s secrets and includes them in a Welcome message directed to Bob, along other contextual information required to participate in the group.
-All other members of the group then process Alice’s commit. Bob’s Key Package also included a leaf node which is inserted into the group’s TreeKEM after validating his credentials.
-External Join
+When the proposal is committed, Alice creates a Welcome message directed to Bob, which contains the information needed to participate in the group.
+All other members of the group then process Alice’s commit.
 In contrast to Add proposals, External Joins are initiated by the joiner, who is not a member of the group.
 The procedure is shown in Figure 1(b): First, the joiner Bob accesses the group’s GroupInfo, which was uploaded to the Directory by the group member Alice.
-Then, he generates a value that will be used to derive the new epoch’s shared secret and encrypts it with the group’s HPKE public key (contained in the GroupInfo) such that it is accessible only to members of the group.
-Finally, Bob generates a commit with the Join proposal - a new proposal type defined by MLS -, which is validated and processed by all members of the group.
-We remark that an MLS group can specify if External Joins are allowed.
-4 Empirical evaluation of MLS
-In this Section we will analyse the most relevant properties of the execution environment in which we will measure the behaviour of MLS using different Delivery Services.
-We will start by defining the parameters that will configure our testbed.
-Then, we will introduce the variables that we will measure, whose values will be predictably affected by the parameter configuration.
-4.1 Parameters
-The most relevant metric for our experiments is group size, that is, the current number of members of the MLS group.
-It is claimed that MLS’ efficiency scales logarithmically with group size [6, 20], and all of our measurements are is significantly influenced by it.
-We will also analyse how the other parameters introduced in this Section interact with group size.
-In addition to group size, we will measure the influence of the following parameters:
-4.1.1 Delivery Service
-The Delivery Service can be instantiated as a centralised server that manages the MLS groups and relays messages to its participants.
-The use of a server help to ensure consistency as it can ensure that all parties receive all messages and in the same order.
-However, the server also represents a single point of failure and may introduce bandwidth limitations, as all messages need to go through a single machine.
-On the other hand, a distributed network of devices that autonomously transmit and forward messages between them can also serve as a Delivery Service.
-In addition to the distribution of messages, the distributed network can also serve as Directory by storing KeyPackages and GroupInfo messages.
-4.1.2 Paradigm
-Users apply modifications to an MLS group by issuing a commit, which updates the shared secret and certain nodes in the Ratchet Tree.
-This can lead to inefficiencies if multiple commits are generated consecutively, as the updated secrets are needlessly overwritten.
-Alternatively, a commit can simultaneously apply multiple modifications through the use of proposals.
-We refer to this two different approaches as paradigms.
-More specifically, we will distinguish between these paradigms:
+Then, he generates a value that will be used to derive the new epoch’s shared secret and includes it in a Join proposal which then committed by Bob and processed by all members of the group.
+4 Testbed Design
+In this Section we describe the scope of our experimental analysis of the performance of MLS.
+We start by introducing the experiments that study different aspects of the evolution of an MLS group.
+Then, we discuss the measurements we will extract from our each of the experiments.
+Our testbed is composed of two elements.
+Firstly, a number of identical emulated MLS clients that independently participate in MLS groups by creating or joining them and issuing and processing updates.
+Secondly, a centralised Publish/Subscribe Delivery Service that forwards messages to group members.
+In this architecture, users can publish messages to different topics such that they are received by all users that are subscribed to said topic.
+Figure 2 shows the handling of topics in the testbed.
+Every group has a different topic to which all current members are subscribed.
+Every user is subscribed to a welcome topic through which it receives all Welcome messages directed to them.
+Additionally, all groups have their own topic that all members are subscribed to, and every handshake and application message of that group is published through it.
+Figure 2: Publish/Subscribe architecture in our testbed.
+As mentioned, our testbed aims to study the empirical performance of MLS under different situations.
+To that end, we have defined a series of experiments, that are characterised by the specific configuration that is shared by all emulated clients.
+In each of them we aim to study the evolution of the group under specific conditions.
+Table 1 shows a summary of the experiments to be performed, which will be detailed in this Section.
+Experiment Batch Paradigm Committer Target users Measurements
+Commit Group Evolution: Paradigm Commit Random 10000 Gen and Proc Time Commit, Welcome and KP Size Latency
+2 Prop Propose-and-Commit Gen and Proc Time Commit Size
+4 Prop
+8 Prop
+External Join External Join
+First Group Evolution: Tree Degradation Commit First Gen and Proc Time (Deep) Commit Size
+Last Last
+100 Users Stable Group Size Random 100 Gen and Proc Time Commit Size
+750 Users 750
+5000 Users 5000
+Table 1: Summary of the performed experiments.
+4.1 Group Evolution
+Our first batch of experiments explores how measurements are affected by the evolution of the MLS group as it grows in size.
+All experiments described in this section will consist on the emulation of an MLS group to which new members are constantly being added.
+The operation to perform and the user that will perform it will depend on the specific experiment.
+4.1.1 Group Updates Paradigm
+In MLS, the following paradigms can be employed to issue updates to the group, including adding new members:
   • 1.
@@
 However, it also introduces an additional overhead: the generation, distribution and processing of proposal messages.
-In the propose-and-commit paradigm, a commit that applies $n$ proposals requires the exchange of $n+1$ messages: each of the $n$ proposal messages and the commit message itself.
-This contrasts with the commit paradigm, in which every handshake message applies exactly one modification.
-We unify these results under the following Average Update cost ($auc$) metric:
-$auc=\frac{cc+\sum_{i=0}^{n}cp_{i}}{n}$
-where $cc$ and $cp_{i}$ are the costs of the commit and its $i$th proposal, respectively.
 We will also consider External Joins in our discussion of paradigms.
-Although External Joins are not incompatible with any of the other two paradigms - as removes and updates can still be either commit or propose-and-commit - they do represent a different method for adding new members to the group that can be compared with them.
-4.1.3 Ratchet Tree State
-Although the TreeKEM protocol aims to achieve to scale logarithmically as the number of users grow, this is only the ideal case: when all the tree branches, included intermediate nodes, are populated.
-These branches can be blanked by some operations, such as invitations of removals under certain conditions [18].
-In the worst case (i.e.
-whenever all intermediate nodes are blanked), the group is structured as a list, which implies that the scaling becomes linear [15].
-The amount and location of blanked nodes depends mostly on which users perform the updates.
-A group reaches its worst possible state if all updates are issued by the same user - the group creator -, as only the nodes in its path to the root will be populated.
-In contrast, the best state is reached when all members issue an update, without considering removes.
-This ensures that all intermediate nodes are populated.
-This can be achieved by only allowing the last user to join to issue updates.
-We will model these scenarios by restricting who is able to issue updates: the First user for the worst state and the Last for the best case.
-We will compare them to a normal scenario in which updates are performed by Random users.
+Although External Joins are not incompatible with any of the other two paradigms —as removes and updates can still be either commit or propose-and-commit— they do represent a different method for adding new members to the group that can be compared with them.
+To study the impact of the chosen paradigm in MLS, we will conduct a total of five experiments that differ on how updates are issued and how new members are inserted into the group: one for the commit paradigm, one for External Joins and three for the propose-and-commit paradigm using different amounts of proposals per commit: 2, 4 and 8.
+We will refer to these experiments as Commit, External Join, 2 Prop, 4 Prop and 8 Prop, respectively.
+4.1.2 Tree Degradation
+(a) A member adds another user, which creates new empty nodes to the tree. As the committer cannot know the private keys of the new nodes, they remain empty.
+(b) A member removes another user. To ensure the removed user cannot decrypt future updates, all of its known nodes are emptied.
+Figure 3: Events that degrade the state of the Ratchet Tree.
+The tree structure of the group state allows MLS to achieve logarithmic efficiency in the best-case scenario.
+However, this only occurs when all the intermediate nodes are populated.
+Figure 3 shows two scenarios in which intermediate nodes may become blanked: during additions to the rightmost side of the tree and removals.
+Blanked intermediate nodes are populated again only when a member whose path to the root includes it issues an update.
+The impact of blanked intermediate nodes is felt on the amount of ciphertext that need to be distributed, as shown in Figure 4.
+(a) Ratchet tree on its optimal state, with all intermediate nodes populated. New commits includes 3 ciphertexts.
+(b) Ratchet tree on its least optimal state, with all intermediate nodes empty. A new commit includes 7 ciphertexts.
+Figure 4: Impact of tree degradation in the performance of MLS for a 8-member group.
+We will analyse the importance of the tree degradation by conducting two different experiments.
+They will measure performance for the worst —all intermediate nodes empty — and best — all intermediate nodes populated— cases.
+The worst possible state is achieved if all updates are issued by the first user to join the group, as shown in Figure 3(a).
+Conversely, the best possible state is reached if every user issues an update; we achieve this by ensuring that only the last member to join can issue updates.
+We refer to this experiments as First (worst possible state) and Last (best possible state).
+4.2 Stable Group Size
+Whereas in our first batch of experiments we analyse how the measurements change as the group evolves in size, the second batch focuses on the performance of group operations at static group sizes.
+Our objective is to measure the variance of the cost of updates at a particular group size.
+Furthermore, we also analyse if this variance remains stable throughout different group sizes, or if larger groups have more volatile computing costs than smaller ones.
+The experiments will unfold as follows: the group evolves normally until it reaches the target group size.
+Then, members will alternate between removing and adding users, such that the group size remains stable.
+We will perform three different experiments for target group sizes 100, 750 and 5000 users, as they represent different orders of magnitude close to a power of 2.
+We refer to this experiments as 100 Users, 750 Users and 5000 Users, respectively.
 4.3 Measurements
-We will measure the following values in our executions:
+In this work we will focus in computational costs, that is, the time it takes to compute certain MLS operations.
+In particular, we will obtain generation and processing times for different types of MLS messages, measured in CPU time.
+Furthermore, we will also analyse the size of MLS message: this includes handshake messages — both commits and proposals — as well as other message types like Welcome and GroupInfo packages.
+Update Cost
+When considering the impact of paradigm, we unify all costs — both time costs and message sizes — to account for the different amount of proposals being applied in a single commit.
+Recall that whereas in the commit paradigm every handshake message applies exactly one modification, in the propose-and-commit paradigm a commit that applies $n$ proposals requires the exchange of $n+1$ messages: each of the $n$ proposal messages and the commit message itself.
+Thus, we define the update cost as
+$\mathsf{UC}=\frac{cc+\sum_{i=0}^{n}cp_{i}}{n},$
+where $cc$ and $cp_{i}$ are the costs of the commit and its $i$th proposal, respectively. Note that in the commit paradigm, $\sum_{i=0}^{n}cp_{i}=0$ and $n=1$, thus the update cost equals $cc$.
+Deep Analysis
+The experiments First and Last respectively represent the least and most optimal states of the Ratchet Tree.
+As shown in Figure 4, this impacts the amount of ciphertexts that must be created to communicate the new state of the group.
+In order to further understand the differences between both experiments, we will additionally perform a deeper analysis for these experiments that subdivides the measurements obtained into categories, each representing a different aspect of the creation of a commit:
+  • 1.
+    Validation: Operations to ensure that the proposals to apply are valid. Of particular interest is the MLS requirement that encryption keys be unique among all members.
+  • 2.
+    Cryptography: TreeKEM operations such as updating and encrypting the intermediate nodes’ secrets and updating the key schedule plus signing and encrypting MLS messages.
+  • 3.
+    Tree: Operations related to parsing the Ratchet Tree such as calculating the path to the root and finding which keys to use to encrypt and decrypt path secrets.
+Additionally, operations that maintain the integrity of the tree, such as and updating the tree hashes and parent hashes every epoch [12].
+  • 4.
+    Welcome: Operations to communicate the state of the group to a new member, such as exporting the GroupInfo package and the Ratchet Tree. Only applicable to Commit generation.
+  • 5.
+    Storage: Operations to provide persistence to the application by saving the group’s state into memory.
 Latency
-In the context of this work, we define an user’s latency with respect to a commit as ”the amount of time between the generation of the commit and the moment in which its changes are applied to the user”.
-Since an MLS group is composed of multiple users, we will measure the average latency and max latency for each commit generated.
-The latter represents the delay between the generation of a commit and the start of the new epoch for all members.
-The time between the generation of a message and its processing is relevant for all messaging applications, but is particularly relevant for handshake messages in MLS: when a member issues a commit, the group remains in an unstable state until all parties have processed the change and thus cannot adequately exchange any other handshake or application messages.
-Commit generation and processing times
-CGKA protocols that employ tree structures, such as MLS, are claimed to have logarithmic complexity in relation to the number of members [6, 20].
-These claims refer to the Communication cost, that is, the amount of information that needs to be exchanged in order to heal the group.
-This metric’s relation to computational costs is unclear, and thus in this work we analyse both generation and processing time.
-We will use CPU time to measure both values.
-Besides group size, we measure how the choice of paradigm and state of the Ratchet Tree affect generation and processing times.
-Message size
-Our analysis includes handshake messages - both commits and proposals, using the Average Update cost metric - as well as other message types like Welcome and GroupInfo packages.
-The latter two include the full Ratchet Tree, so their size is expected to scale poorly as the group grows.
-This experimental measurement complements the communication cost that has been widely analysed in the literature.
-Our measurements can provide specific values that can be fed into communication cost calculations such that the obtained results can be applied to real-world scenarios.
-5 Implementation
-In this Section we will introduce our implementation of the different Delivery Services we will employ in our execution environment.
-We have developed our implementation in Rust, using the OpenMLS project [29] as a baseline and incrementally expanding upon it.
-We employ two different Delivery Services, each of them with their corresponding Directory.
-The first employs a centralised server and forwards messages through the MQTT protocol while the second distributes them through a peer-to-peer network.
-Their implementations are analysed in subsequent sections.
-Both are based in a Publish-Subscribe architecture.
-Figure 3 shows the handling of topics of a Publish/Subscribe Delivery Service for MLS.
-Every group has a different topic to which all current members are subscribed.
-Every user is subscribed to a welcome topic through which it receives all Welcome messages directed to them.
-Additionally, all groups have their own topic that all members are subscribed to, and every handshake and application message of that group is published through it.
-Figure 3: Subscription process for new group members.
-5.1 MQTT Delivery Service
+Finally, we also measure latency, defined as the time between the instant a message is sent by a committer and the instant it is received by a processer.
+Whenever a commit is issued but has not been processed yet by all members, the group is in an inconsistent state as members disagree on the current epoch and group state: latency determines how long this inconsistency may last.
+To avoid redundancy, latency does not include neither commit generation nor processing times.
+5 Emulation Environment
+In this Section we introduce the implementation of the presented testbed.
+Our code [^2] Available at https://github.com/SDABIS/mls_experimental_analysis consists mostly of Rust code that implements the emulated MLS client, as well as the Docker infrastructure required to execute the Delivery Service and multiple clients.
+5.1 Delivery Service
 MQTT is a protocol for establishing publish/subscribe queues of messages [28].
-In this architecture, users can publish messages to different topics such that they are received all users that are subscriber to said topic.
-The transmission of these messages to all subscribers is managed by a centralised server named MQTT broker.
+The transmission of these messages to all subscribers is managed by a MQTT broker.
 For our execution environment, we employ the open-source Mosquitto [27] implementation of a MQTT broker.
@@
 The latter needs to be updated for each new group epoch, so users aiming to perform an External Join can successfully access the group.
-Furthermore, this web server also acts as a signaling server in which users can register to notify other participants of their existence, such that they can be invited to MLS groups through an Add proposal.
-5.2 GossipSub Delivery Service
-For this Delivery Service we maintain the Publish-Subscribe nature but replace the centralised MQTT server with a distributed architecture.
-We employ the GossipSub protocol [32] that allows peers to subscribe to different topics; peers share with each other metadata of messages they have received and can request the full message if it is from a topic they are subscribed to.
-Connections are dynamic and favour other members subscribed to the same topic.
-We also employ a Decentralised Hash Table (DHT) as this Delivery Service’s Directory.
-Each client publishes its KeyPackage as a key-value pair, with the key referencing its user identifier.
-We limit the involvement of the server to that of signaling server as defined for the MQTT Delivery Service.
-For our implementation we include Rust’s Libp2p [33] into the simulated clients, which provides an implementation for GossipSub and the Kademlia [34] DHT.
-5.3 Client Simulation
-Our testbed is composed of multiple simulated clients that independently participate in MLS groups.
+Furthermore, this web server also acts as a signaling server in which users can register to notify other participants of their existence, such that they can be invited to MLS groups.
+5.2 Client Emulation
+Our testbed is composed of multiple emulated clients that independently participate in MLS groups.
 They emulate the behaviour of real clients by randomly adding or removing members, updating their individual leaf nodes and sending application messages.
-[cgka] ds = "mqtt" groups = ["group_1", "group_2"] external_join = true join_chance = 0.01 issue_update_chance = 0.2 message_chance = 0.3 scale = false auth_policy = "Random" message_length_min = 500 message_length_max = 2000 sleep_millis_min = 20000 sleep_millis_max = 60000 [paradigm] paradigm = "propose" proposals_per_commit = 4 invite_chance = 0.6 remove_chance = 0.1 update_chance = 0.3 [http_server] url = "http://<ip>:<port>" [mqtt] url = "tcp://<ip>:<port>" [meta] replicas = 10
-Figure 4: Example Client configuration file.
-The behaviour of our simulated clients can be specified in detail through configuration files, as shown in Figure 4.
-Among other parameters, it is possible to configure the groups the user will try to join (or create if they do not exist), how often it will attempt to perform an action and with what probability.
-The configuration file also defines which paradigm to use and the types of actions (Add, Remove, Update, Join) the client is execute, as well as the Delivery Service that will be employed.
-group_1 8 User_0d4341e0c0f4 Invite User_b3f20ed42c2a 1065 1739176120380282661 5885 group_1 8 User_d13041578e84 Process User_0d4341e0c0f4 1739176120384444036 8599 group_1 8 User_b3f20ed42c2a Welcome User_0d4341e0c0f4 2755 1739176120461981624 1823
-Figure 5: Logs generated by clients.
-Every action related to the evolution of an MLS group is recorded in a log file.
-Figure 5 shows the trace of an Invitation, which includes involves the following entries: Invite, Process and Welcome.
-Each entry records the size of the message, the timestamp at which it was performed and its time cost.
+Clients interact with each other through real MLS messages sent through the Delivery Service.
+We have developed our implementation of an emulated client in Rust, using the OpenMLS project [31] as a baseline and incrementally expanding upon it.
+Clients can be deployed in Docker containers, each executing a configurable amount of copies.
+The behaviour of Clients can be extensively configured to reach specific group states.
+Each client has a different name and a list of groups they aim to join.
+There are two behaviours clients can follow to decide when to act in a group.
+In the independent mode, clients randomly decide whether or not to send a message at fixed time intervals.
+This may result in redundant messages if multiple users attempt to commit at the same time.
+In orchestrated mode, the transcript hash of the group is used as a seed to a RNG, such that all group members agree on a randomly chosen client who issues the next update.
+This allows for better coordination resulting in faster executions.
+In either mode, clients always process incoming messages immediately after their arrival.
+Clients can also be configured to only consider issuing updates if they are the first or last member to join the group, as required by the Tree Degradation experiments described in Section 4.
+When they are tasked with issuing updates, clients will follow one of the paradigms described in Section 4.
+If the propose-and-commit paradigm is employed, clients will issue proposals until a threshold of proposals is reached and then commit them.
+Clients may also perform an External Join to a group in which they are not members.
+Clients randomly choose between adds, updates and removals according to configured proportions.
+In order to reach groups of stable size for the experiments introduced in Section 4, this proportions can be overridden once the group reaches a threshold of members such that only removes may be issued.
+Additionally, clients may send randomly-generated application messages.
+The emulated clients record every action they perform in detailed log file that includes the timestamp and duration of the action.
 6 Experimental Results
 In this section we deploy the environment presented in previous sections and perform empirical evaluations on its performance.
-We have executed all following tests three times with the objective of clearing any potential noise in the measurements.
-Every figure in this Section shows the average of the results obtained in all executions; we omit the standard deviation as it is not relevant to our scenarios.
-6.1 Latency
-(a) MQTT
-(b) GossipSub
-Figure 6: Mean and Max Latency for the Delivery Services.
-Figures 6(a) and 6(b) show the latency for the two Delivery Services.
-While both experience a significant increase as the number of users grow, the GossipSub DS is noticeably less efficient as its latency more than doubles that of the MQTT DS.
-Latency also increases more irregularly of GossipSub, with an important increase at 512 group members.
-However, at higher group sizes the latency grows slower for GossipSub.
-6.2 Paradigm
-We now compare the Average Update cost for the commit and propose-and-commit paradigms, including multiple amounts of proposals per commit for the latter to see how the cost varies as the number of updates increases.
+All experiments in Table1 have been executed by providing their respective configuration to the emulated clients.
+The target group size for all experiments —except those related to stable group size— is 10000.
+We note that in all experiments clients employ the orchestrated behaviour since it allows for faster and more controlled executions.
+All experiments were executed on an Ubuntu 20.04 virtual machine equipped with 128 GB of RAM and 16 CPU cores operating at 3.2 GHz.
+The 10,000 emulated clients were divided into 10 Docker containers each running 1000 instances.
+The Delivery Service —including both the HTTP Directory and the MQTT Broker— were run in a similar virtual machine located in the same local network.
+6.1 Group Evolution
+In order to ensure groups grew in size, clients were configured with probabilities of 60% invite, 20% update and 20% remove.
+Every experiment has been executed 3 times; the data shown in this Section are the aggregation of the results obtained in every execution.
+For each of the following graphs, the 10,000 measurement points have been subsequently aggregated into 100 points, with each point representing the aggregation of 100 measurements.
+Figure 5 shows the mean and maximum latency in the experiment Commit. Both measurements increase slowly as the group grows and the consumed bandwidth increases.
+Figure 5: Mean and Maximum Latency for the Delivery Service as the number of group members grows.
+6.1.1 Paradigm
+We now compare the update cost in our Paradigm experiments.
 We note that our measurements do not include the cost of retrieving information from the Directory such as KeyPackages for Add proposals, as it remains constant regardless of the chosen paradigm.
-(a) Commit and propose-and-commit paradigms
-(b) Commits (also including Welcome processing time) and External Joins.
-Figure 7: Average Update generation time for different paradigms as the number of users grow.
+(a) Experiments Commit, 2 Prop, 4 Prop and 8 Prop.
+(b) Experiment External Join, with Commit shown for comparison. Welcome also including Welcome processing time — and External Joins.
+Figure 6: Update generation time for different paradigms as the number of group members grow. Update cost is calculated by dividing total cost by number of proposals applied in a single commit.
+Figures 6(a) shows the generation time for the experiments Commit, 2 Prop, 4 Prop and 8 Prop, whereas Figure 6(b) shows the same measure for the experiment External Join.
+Clearly, the update cost is significantly reduced as the number of proposals increases.
+We highlight that an user performing an External Join must also parse the current state of the group including the full Ratchet Tree, which would explain the poor scaling of External Joins.
+This operation is also performed in an Invitation setting by the invited user when the Welcome message is processed; when accounting for the processing time of Welcome messages, the two methods have a similar cost (as shown in Figure 6(b) by the line labelled Commit + Welcome).
 (a) Processing time.
 (b) Message size.
-Figure 8: Average Update cost for different paradigms as the number of users grow.
-Figures 7(a) and 7(b) show the Average Update generation time for the two paradigms and for External Joins, respectively.
-Clearly, the Average Update cost is significantly reduced as the number of proposals increases.
-We also note that the time to perform an External Join is much larger to that of inviting new members through Add proposals.
-In order to understand these differences we highlight that an user performing an External Join must also parse the current state of the group including the full Ratchet Tree, which would explain the poor scaling of External Joins.
-This operation is also performed in an Invitation setting by the invited user when the Welcome message is processed; when accounting for the processing time of Welcome messages, the two methods have a similar cost (as shown in Figure 7(b) by the line labelled Commit + Welcome).
+Figure 7: Update cost for different paradigms as the number of users grow. Update cost is calculated by dividing total cost by number of updates of proposals applied in a single commit.
 Figure 7(a) shows the processing time per Update as the number of users grow.
 As with generation time, the processing cost is significantly reduced as the number of proposals increases.
-Processing time is noticeably more stable than generation time, although it experiences significant increases at the powers of 2 as a new layer is added to the Ratchet Tree.
-It is apparent that both generation and processing times scale mostly linearly with the number of members in the CGKA group, with $R^{2}$ scores over 0.9 when adjusting for linear regression and around 0.6 for logarithmic regression.
 Similarly, Figure 7(b) shows the message size per update.
 Clearly, the correlation between the message size and the number of users is much smaller than in the other metrics.
-Contrary to the results obtained while measuring processing and generation times, message size is generally bigger as the number of proposals increase.
-We recall that commit messages also include copies of the proposals it applies; this redundant information increases the cost per update.
-6.3 Other Messages
-Figure 9: Message size for Welcome messages and GroupInfo packages.
+The increases at powers of 2 is also noticeable in message size, as shown by the spikes at 4096 and 8192 members.
 We also evaluate the size of the Welcome messages and GroupInfo packages, which take part in Invitations and External Joins, respectively.
@@
 Since both contain a full copy of the Ratchet Tree, they are significantly larger than any handshake message.
-Expectedly, their size increases linearly with the number of group members.
-6.4 Ratchet Tree State
-(a) Generation time.
-(b) Message size.
-Figure 10: Generation time and message size for different states of the group’s Ratchet Tree. Processing time is omitted as it is not influenced by this metric.
-Finally, we analyse the impact in performance of the current state if the Ratchet Tree, exemplified by the First, Last and Random scenarios introduced in Section 4.1.
-Figure 10 shows the generation time and message size - processing time is omitted as no significant difference was found between the three scenarios.
-As with previous measurements, generation time also scales linearly in all scenarios - this includes Last which corresponds to the best possible state of the tree.
-Message size does show a significant correlation, with First increasing much faster than in the other scenarios.
+Figure 8: Message size in Bytes for Welcome messages and GroupInfo packages.
+6.1.2 Tree Degradation
+Figure 9: Message size for the First and Last experiments as the number of users grow. The results of the Commit experiment are included for comparison.
+Figure 9 shows the message size for the experiments First and Last.
+Since in the former almost all intermediate nodes are blanked, the commit messages must contain one ciphertext for every member and thus their size grows linearly.
+(a) Experiment First.
+(b) Experiment Last.
+Figure 10: Generation time measured for the experiments First and Last that respectively correspond to the worst and best possible Ratchet Tree states.
+Measures are subdivided into categories.
+Storage is cropped to improve visibility, as it increases linearly and experiences significant jumps at powers of two.
+Figure 10 shows the distribution of time cost across different categories for commit generation in both experiments.
+Whereas most categories predictably exhibit the same behaviour, Cryptography is significantly more costly in First than in Last, as the path secrets need to be encrypted once for every member.
+In both, Processing time is omitted as the cost distribution across categories was identical to that of commit generation.
+Notably, Cryptography represents a small percentage of the total cost: at 10,000 users, it reaches 1% for the experiment Last and 10% for First.
+6.2 Stable group size
+In this Section we execute the experiments related to stable group size: 100 Users, 750 Users and 5000 Users.
+For each experiment we measure 10,000 commits at the corresponding group size.
+Figure 11 shows the measured commit generation and processing time and message size.
+The latter metric experiences a noticeably higher variance than the others.
+In particular, commit processing is the more stable metric.
+All of them follow a normal distribution, which is more easily observable for larger groups.
+(a) Commit size.
+(b) Generation time.
+(c) Processing time.
+Figure 11: Commit size and generation and processing time for groups of 100 (top), 750 (middle) and 5000 (bottom) members.
+The dotted and dashed lines respectively represent the median and mean of each series.
 7 Discussion
-Comparison with other works
-Protocol Perspective Group Size Time Bandwidth Tree State Paradigm Join DS
-[6] CGKA Variant Theoretical N/A
-[3, 4] CGKA Theoretical N/A
-[11] CGKA Variant Experimental $\approx$ 128
-[10] CGKA Variant Experimental $\approx$ 128
-[24] MLS Theoretical N/A
-Ours MLS Experimental $\approx$ 1000
+In this Section we analyse the data introduced in the previous Section and discuss its real-world implications.
+We also compare our work to other analysis of performance of MLS or other CGKA variants.
+Finally, we discuss the limitations encountered and propose future work.
+7.1 Analysis
+Message cost.
+Generation and processing times are comparable and represent a small cost in the order of milliseconds.
+The propose-and-commit paradigm also significantly affects performance: the cost of generating and processing proposal messages is negligible compared to the cost of applying an update to the group.
+As shown, more proposals per update increase the efficiency of the protocol.
+The size of handshake messages is less affected by the amount of group members and the paradigm, as it is mostly determined by the current state of the Ratchet Tree.
+External Joins combine the operations of generating a commit and processing a Welcome message, as the joiner needs to parse the full Ratchet Tree in order to join.
+When taking these two operations into account, the cost of an Invitation is similar to that of an External Join.
+As shown in Figure 11, commit messages at a given group size can have significantly different sizes.
+This is determined by the number of ciphertexts contained in them, which in turn depends on the number of blanked intermediate nodes as explained in Figure 4.
+Notably, commit generation and processing times have less variance as their cost is mostly determined by other operations unrelated to the cryptographic component of MLS.
+The increased bandwidth of larger groups has a noticeable impact on latency, as the cost of distributing messages between all members increases.
+However, this metric grows slower than commit generation and processing times.
+When comparing Figure 5 with Figures 6 and 7, mean latency is around 2.5 times more expensive than commit generation and processing at 1000 members, whereas their cost is comparable at 10,000 members.
+Ratchet Tree Size.
+Our results show that the overwhelming majority of the costs both in generation and processing are caused by the size of the full Ratchet Tree.
+Indeed, Figure 10’s categories Storage, Welcome, Validation and Tree all scale linearly with the size of the group’s state; only Cryptography is independent, but its cost is negligible in comparison.
+The result is that the overall scaling of MLS’s performance is linear, which contradicts the theoretical results which point to logarithmic performance [10, 9, 16].
+We remark that Storage cannot be considered optional in practical scenarios, as it is required to provide persistence between sessions.
+Even then, the remaining categories still scale linearly and their cost cannot be circumvented: Tree includes the calculation of tree hashes and parent hashes that prevent insider attacks [8] and Validation’s check that all keys should be different is needed to ensure Post-Compromise Security.
+Furthermore, the most expensive operations of the protocol are processing a Welcome and performing an External Commit, as shown in Figure 6(b).
+Both require processing the full Ratchet Tree, whether it be from a Welcome message or a GroupInfo package.
+Thus, the most relevant factor to consider regarding performance in MLS is the size of the Ratchet Tree, which scales linearly by the number of members.
+The logarithmic complexity of the cryptographic operations only applies to a negligible step of the computation.
+7.2 Real-world Implications
+Figure 5 shows that maximum latency can reach up to 0.3 seconds, which could potentially be much higher depending on the network or Delivery Service employed.
+This measure determines the amount of time in which the group is in an inconsistent state, since some members have not yet joined the current epoch.
+Whereas for most applications this span is too short to be noticed, it is something that must be taken into account for situations that require strict synchronisation among group members.
+The results of the experiment First show that groups in which only a small number of users generate updates are less efficient, with a particularly high commit size.
+This is significant as real-world messaging applications usually limit updates to one or more administrators.
+Thus, members should be allowed to periodically update their own leaf node to prevent the accumulation of blanked intermediate nodes.
+As shown in Figure 10, the size of the Ratchet Tree has a significant impact in performance as this structure needs to be managed to maintain its integrity and stored in memory at every commit.
+This means elements that appear unrelated to performance but increase the size of the tree (e.g.
+clients using large credentials or signature keys) may provoke an unexpected increase in the cost of group operations.
+7.3 Comparison with other works
+Table 2 shows a comparison between our work and other works that perform an analysis of performance of CGKA.
+Our experimental approach allows us to measure CPU time in generating and processing messages.
+Our work covers topics whose impact in performance has not been studied in any other of the listed works, such as External Joins, the choice of paradigm and the importance of the Delivery Service.
+Notably, other works that analyse the theoretical performance of CGKA [10, 9, 4, 17] focus on communication cost, which is defined as the number of ciphertexts required to heal from corruption.
+Our experimental analysis serves as a complement to these theoretical bounds by providing a real-world reference of the size of the exchanged messages as well as the cost of generating and processing them.
+Furthermore, a theoretical approach does not adequately capture the cost of maintaining and storing the Ratchet Tree, which our analysis shows to be much higher than that of cryptographic functions.
+Work Protocol Perspective Group Size Time Bandwidth Tree State Paradigm Join DS
+Alwen et al. [7] Variant (SAIK) Theoretical N/A
+Auerbach et al. [10] CGKA Theoretical N/A
+Anastos et al. [9] CGKA Theoretical N/A
+Weidner et al. [34] Variant (DGKA) Experimental $\approx$ 128
+Chevalier et al. [18] MLS Experimental $\approx$ 1,000
+Balbás et al. [11] Variant (A-CGKA) Experimental $\approx$ 128
+Paillat et al. [30] MLS Theoretical N/A
+Ours MLS Experimental $\approx$ 10,000
 Table 2: Comparison between works that analyse CGKA performance and ours.
 The columns indicate if the works consider the following topics.
-CPU: update generation/processing time.
+Time: update generation/processing time.
 Bandwidth: message size and communication cost.
+Tree State: Influence of the current state of the Ratchet Tree.
 Paradigm: influence of chosen paradigm.
-Tree State: Influence of the current state of the Ratchet Tree.
 Join: External Joins.
 DS: influence of the Delivery Service.
-Table 1 shows a comparison between our work and other works that perform an analysis of performance of CGKA.
-Our experimental approach allows us to measure CPU time in generating and processing messages and the importance of the Delivery Service in their distribution.
-When compared with other experimental works, ours has the largest group size and also covers less studied topics such as the influence of paradigm and the current state of the Ratchet Tree.
-Delivery Service
-The increased bandwidth of larger groups has a noticeable impact on latency, as the cost of distributing messages between all members increases.
-The MQTT DS performs significantly better than the GossipSub DS, but the former does not require a centralised server to distribute the messages and thus is more reliable.
-Latency represents the most significant cost in the evolution of MLS groups, as every other generation or processing cost is negligible in comparison.
-Updates in MLS are necessarily sequential: if two commits are generated from the same epoch, the group enters into an inconsistent state.
-In order to generate a valid commit, an user must have received and processed all previous valid commits.
-Thus, latency is a significant bottleneck to group evolution as it represents the minimum amount of time between two valid commits.
-Our measurements reach up to 2 seconds of maximum latency in which the group is in an inconsistent state; while for most applications this span is too short to be noticed, it is something that must be taken into account for situations that require strict synchronisation.
-Message cost
-Generation and processing times are comparable and represent a small cost in the order of milliseconds.
-Both experience noticeable increases when a new layer is added to the Ratchet Tree, that is, when group size reaches the next power of 2.
-In practice, both generation and processing times increase linearly with the amount of users in the group, which contradicts the theoretical assumptions that the protocol complexity is logarithmic.
-Notably, this trend is maintained even in the best-case scenario where the minimum amount of intermediate nodes are blanked.
-When compared to latency, with generation and processing times represent a negligible cost.
-Welcome messages and GroupInfo packages are significantly costlier to process - as shown in Figure 7(b) - than normal commits.
-Fortunately, they only need to be processed once in order to access a group.
-This cost difference is due to their size as they both contain the full Ratchet Tree, which needs to be parsed by the receiving party.
-In contrast, the size of handshake messages is less influenced by group size and is mostly determined by the state of the Ratchet Tree.
-Generation time is also significantly affected.
-Even though these two metrics are slightly correlated they present significant differences and thus should be analysed separately.
-Our results show that groups in which only a small number of users generate updates are less efficient.
-Impact of paradigm
-The propose-and-commit paradigm also significantly affects performance: the cost of generating and processing proposal messages is negligible compared to the cost of applying an update to the group.
-As shown, more proposals per update increase the efficiency of the protocol, but we remark that such amount of modifications to a group applied at once would be rare in a real setting, only possible in a highly unstable group.
-The size of handshake messages is less affected by the amount of group members and the paradigm, as it is mostly determined by the current state of the Ratchet Tree.
-External Joins combine the operations of generating a commit and processing a Welcome message, in the sense that the joiner needs to parse the full Ratchet Tree in order to join.
-When taking these two operations into account, the cost of an Invitation is similar to that of an External Join.
+When compared with other experimental evaluations such as [34, 18, 11], the increased depth of our analysis allows us to extract more robust conclusions.
+Indeed, the increments in computing cost at group sizes that are powers of 2 are not captured by works that limit this metric to 100 or 1,000 members.
+Our analysis also shows that cryptographic operations represent a very small percentage of experimental cost.
+Thus, claims about a CGKA variant’s efficiency must take into account the disproportionate cost of storage and tree management that may be unrelated to its contribution.
+7.4 Limitations
+In this work, we employ an ideal Delivery Service that is readily accessible to all members and that ensures that messages are delivered in order.
+Thus, our analysis does not capture real-world inconveniences that can be introduced by the network such as limited bandwidth, packet losses or out-of-order delivery.
+MLS has also sparked interest for its application in distributed environments [3, 32, 26]; replacing the centralised Delivery Service employed in our testbed with a distributed one could be of interest to analyse its implications on performance.
+Our analysis is limited to 10,000 users because of the significant cost of the execution of the experiments.
+For reference, a single commit in a 10,000-member group takes 0.1 seconds to be processed: if executed sequentially, it takes more than 15 minutes for all members to process it.
+In order to measure performance in larger groups, a stronger effort in parallelism is required.
+Alternatively, larger analysis could also ignore memory storage in order to remove the largest computational cost.
+The results obtained in this work show the relevance of handling the state of the Ratchet Tree.
+Thus, future work could focus on studying alternatives that minimise its size, such as compression.
+Similarly, the performance of clients that do not possess a copy of the Ratchet Tree —such as the recently proposed Partial MLS [24]— could be compared to our results.
 8 Conclusion and Future Work
@@
 We analyse the protocol under different parameters such as Delivery Service, chosen paradigm or state of the Ratchet Tree.
-Our results show that theoretical claims of logarithmic complexity in communication cost do not necessarily manifest in practice.
-Our work also demonstrates the relevance of practical considerations in the performance of MLS, as opposed to theoretical analysis.
-We have found that the most significant time cost - by multiple orders of magnitude - is message distribution through the Delivery Service.
-Furthermore, the messages that introduce higher workload are Welcome and GroupInfo messages, as they contain the full Ratchet Tree.
-To the extent of our knowledge, no other work has attempted to increase the efficiency of these operations.
-We also publish our execution environment as open source, which includes a configurable simulated MLS client and two different Delivery Services.
-We hope that it can help other researchers to perform their own simulations and that it can serve as a baseline for more complex simulation environments.
-We plan on continuing to study the applicability of concepts defined specifically in MLS - such as the propose-and-commit paradigm or External Joins [13] - to general definitions of CGKA and their impact on the security properties and efficiency of these schemes.
-Declaration of competing interest
-The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
-Acknowledgement
-D.S. acknowledges support from Xunta de Galicia and the European Union (European Social Fund - ESF) scholarship [ED481A-2023-219].
-This work is part of the project TED2021-130369B-C31 and TED2021-130492B-C21 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”.
+Our results show that theoretical claims of logarithmic complexity in communication cost do not manifest in practice.
+Our work also demonstrates the relevance of practical considerations in the performance of MLS.
+We have found that the most significant time costs are message distribution through the Delivery Service and the storage of the Ratchet tree, which are not taken into account in theoretical analysis.
+We also publish our execution environment as open-source, which includes a configurable emulated MLS client and two different Delivery Services.
+Our testbed can be employed by other researchers to perform in-depth studies of the behaviour of MLS in different environments, and can serve as a baseline for the development of more complex experiments.
+Our implementation can be easily adapted to execute other CGKA variants, such that their performance can be systematically compared to that of MLS.
+Future work should focus on implementing a more complex Delivery Service that can accurately represent real-world inconveniences and optimising the emulated client’s memory usage and parallelism to allow for larger experiments.
+Acknowledgements
+The work is funded by the Plan Complementario de Comunicaciones Cuánticas, Spanish Ministry of Science and Innovation(MICINN), Plan de Recuperación NextGeneration, European Union (PRTR-C17.I1, CITIC Ref.305.2022), and Regional Government of Galicia (Agencia Gallega de Innovación, GAIN, CITIC Ref.
+306.2022).
+D.S.
+acknowledges support from Xunta de Galicia and the European Union (European Social Fund - ESF) scholarship [ED481A-2023-219].
+We also acknowledge support from the Xunta de Galicia and the European Union (FEDER Galicia 2021-2027 Program) Ref.
+ED431B 2024/21, CITIC ED431G 2023/0.